# Stability of ranks under field extensions

### v3

@@ 3. Rank and subrank of a field extension
$N>g_{m}-1>\frac{g_{m+1}}{2q}-1>\frac{n-3}{4qd}-1\geq\frac{n}{12qd}.$
For $n<2dg_{3}+3\leq 2dq^{3}$, we split the discussion into two cases. If $q^{2}\geq n/d$, Proposition 3.1 implies that $\sigma_{q^{2}}(n)\geq n/d$. If $q^{2}<n/d$, a similar interpolation argument as in the proof of Proposition 3.1 shows $\sigma_{q^{2}}(n)\geq q^{2}$. Therefore, we obtain $\sigma_{q^{2}}(n)\geq\min(n/d,q^{2})\geq\frac{1}{2qd}n$. To complete the proof, we recall from Lemma 3.6 that
$\sigma_{q}(2n)\geq\sigma_{q}(2)\sigma_{q^{2}}(n)\geq\frac{n}{12qd}.$
@@
Lemma 3.5 implies $\sigma_{q}(l)\geq\sigma_{q}(2n)\geq n/12qd>l/48qd^{2}$ and this completes the proof. ∎
4. Analytic rank and Geometric rank under field extension
@@
This section is devoted to a discussion of the behaviour of $\operatorname{AR}_{\mathbb{K}}(T)$ and $\operatorname{GR}_{\mathbb{K}}(T)$ under field extensions, defined respectively in (6) and (4). For each $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and a field extension $\mathbb{F}/\mathbb{F}_{q}$, we denote by $T^{\mathbb{F}}$ the tensor in $\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ determined by $T$. Let $T_{d,\mathbb{F}}$ be the tensor in $\mathbb{F}^{\otimes d}$ corresponding to the multiplication map $\varphi_{d}$ defined in (7). By definition, $T\boxtimes T_{d,\mathbb{F}}$ is a tensor over $\mathbb{F}_{q}$ and we have
$T^{\mathbb{F}}=T\boxtimes T_{d,\mathbb{F}}.$
Here $T^{\mathbb{F}}$ is regarded as a tensor over $\mathbb{F}_{q}$. Moreover, we remark that $T^{\mathbb{F}}\neq T$ as tensors over $\mathbb{F}_{q}$.
Lemma 4.1.
Given $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and $S\in\mathbb{F}_{q}^{m_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{m_{d}}$, if there exist $M_{i}\in\mathbb{F}_{q}^{n_{i}\times m_{i}}$ for each $1\leq i\leq d$ such that $(M_{1}\otimes\cdots\otimes M_{d})S=T$, then $\operatorname{AR}_{\mathbb{F}_{q}}(T)\leq\operatorname{AR}_{\mathbb{F}_{q}}(S)$.
Proof.
Since $M_{1}\otimes\cdots\otimes M_{d}=(M_{1}\otimes I_{m_{2}}\otimes\cdots\otimes I_{m_{d}})(I_{n_{1}}\otimes M_{2}\otimes\cdots\otimes I_{m_{d}})\cdots(I_{n_{1}}\otimes I_{n_{2}}\otimes\cdots\otimes M_{d})$, it suffices to assume that $m_{i}=n_{i}$ and $M_{i}=I_{n_{i}}$ for each $11\leq i\leq d-$ 1. Denote $N\coloneqq\prod_{i=1}^{d-1}n_{i}$. Let $T^{(d)}\in\mathbb{F}_{q}^{n_{d}\times N}$ (resp. $S^{(d)}\in\mathbb{F}_{q}^{m_{d}\times N}$) be the $d$-th flattening of $T$ (resp. $S$). Then we have
$(M_{d}\otimes I_{N})S^{(d)}=T^{(d)}.$
In particular, $Z_{d}\left(S\right)\subseteq Z_{d}\left(T\right)$. By Lemma 2.2, we have
@@
$\operatorname{AR}_{\mathbb{F}_{q}}(T)$ $=\sum_{i=1}^{d}n_{i}-n_{d}-\log_{q}\left(\left\lvert Z_{d}\left(T\right)\right\rvert\right)$
$\leq\sum_{i=1}^{d}n_{i}-n_{d}-\log_{q}\left(\left\lvert Z_{d}\left(S\right)\right\rvert\right)$
$=\operatorname{AR}_{\mathbb{F}_{q}}(S).\qed$
@@
$\operatorname{AR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right)$ $=\sum_{i=1}^{d}n_{i}-n_{d}-\log_{\lvert\mathbb{F}\rvert}\left(\left\lvert Z_{d}\left(T^{\mathbb{F}}\right)\right\rvert\right)$
$=\sum_{i=1}^{d}n_{i}-n_{d}-\frac{\log_{|\mathbb{F}_{q}|}(\lvert Z_{d}(T\boxtimes T_{d,\mathbb{F}})\rvert)}{l}.$
Since $\mathbb{F}_{q}^{n_{i}}\otimes\mathbb{F}\simeq\mathbb{F}_{q}^{n_{i}l}$ as $\mathbb{F}_{q}$-vector spaces, we may regard $T\boxtimes T_{d,\mathbb{F}}$ as a tensor in $\mathbb{F}_{q}^{n_{1}l}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}l}$. By Lemma 2.2 again, we have
$\log_{\lvert\mathbb{F}_{q}\rvert}\left(\lvert Z_{d}(T\boxtimes T_{d,\mathbb{F}})\rvert\right)=\sum_{i=1}^{d}n_{i}l-n_{d}l-\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes T_{d,\mathbb{F}}).$
@@
$\operatorname{AR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right)=\frac{\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes T_{d,\mathbb{F}})}{l}.$ (13)
Denote $s\coloneqq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$. By definition, there exist matrices $M_{i}\in\mathbb{F}_{q}^{s\times l}$ for $1\leq i\leq d$ such that $(M_{1}\otimes\cdots\otimes M_{d})T_{d,\mathbb{F}}=\operatorname{Id}_{s}$, where $\operatorname{Id}_{s}$ is the diagonal tensor in $(\mathbb{F}_{q}^{s})^{\otimes d}$. Therefore, we have
$\left((I_{n_{1}}\boxtimes M_{1})\otimes\cdots\otimes(I_{n_{d}}\boxtimes M_{d})\right)(T\boxtimes T_{d,\mathbb{F}})=T\boxtimes\operatorname{Id}_{s}=T^{\oplus s}.$
@@
$\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes_{\mathbb{F}_{q}}T_{d,\mathbb{F}})\geq\operatorname{AR}_{\mathbb{F}_{q}}(T^{\oplus s})=s\operatorname{AR}_{\mathbb{F}_{q}}(T)\geq c_{d,q}l\operatorname{AR}_{\mathbb{F}_{q}}(T)$ (14)
@@
To conclude the proof, we replace $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$ by $\operatorname{R}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$ in the above argument and the other inequality follows immediately from Theorem 3.2. ∎
@@
It is proved in [28, Theorem 8.1] that for any $T\in\mathbb{Z}^{n_{1}}\otimes\cdots\otimes\mathbb{Z}^{n_{d}}$,
$\liminf_{p}\operatorname{AR}_{\mathbb{Z}_{p}}\left(T^{\mathbb{Z}_{p}}\right)=\operatorname{GR}_{\mathbb{Z}}(T),$ (17)
@@
According to [1, Proof of Theorem 1.13 (1)], we have $C\leq(1-\log_{q}(d))^{-1}$. Furthermore, the proof of Theorem 4.2 indicates that $c\geq(48d^{2}q)^{-1}$. As far as we are aware, the existence of $c$ is not previously known. Let $\mathbb{F}$ be a field and let $\langle n,n,n\rangle_{\mathbb{F}}$ be the tensor in $\mathbb{F}^{n}\otimes\mathbb{F}^{n}\otimes\mathbb{F}^{n}$ corresponding to the $n\times n$ matrix multiplication. By [28, Theorem 3], we have $\operatorname{GR}_{\mathbb{F}}(\langle n,n,n\rangle)=\lceil 3n^{2}/4\rceil$. Applying Theorems 4.2, 4.3 and Proposition 4.4 to $\langle n,n,n\rangle_{\mathbb{F}_{q}}$, we obtain the following result, which may be of independent interest.
@@
5. Partition rank and slice rank under field extension
In this section, we investigate the behaviour of the partition rank and the slice rank under field extensions. Let $\mathbb{K}$ be a field and let $\mathbb{F}$ be a finite extension of $\mathbb{K}$. We notice that any tensor $S\in\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ can be viewed as a tensor over $\mathbb{K}$ via the $\mathbb{K}$-linear isomorphism:
$\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}\simeq\left(\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}\right)\otimes_{\mathbb{K}}\mathbb{F}.$
Thus, both $\operatorname{PR}_{\mathbb{F}}(S)$ and $\operatorname{PR}_{\mathbb{K}}(S)$ are defined. Moreover, they are related by an inequality.

### v4

@@ 3 Rank and subrank of a field extension
$N>g_{m}-1>\frac{g_{m+1}}{2q}-1>\frac{n-3}{4qd}-1\geq\frac{n}{24qd}.$
The last inequality can be verified as follows. This inequality is equivalent to $5n\geq 18+24qd$. Since $n\geq 2dg_{3}+3=2d(q^{3}-2q+1)+3\geq 5dq+3$, we have $5n\geq 5(5dq+3)\geq 24dq+18$. For $n<2dg_{3}+3\leq 2dq^{3}$, we split the discussion into two cases. If $q^{2}\geq n/d$, Proposition 3.1 implies that $\sigma_{q^{2}}(n)\geq n/d$. If $q^{2}<n/d$, a similar interpolation argument as in the proof of Proposition 3.1 shows $\sigma_{q^{2}}(n)\geq q^{2}$. Therefore, we obtain $\sigma_{q^{2}}(n)\geq\min(n/d,q^{2})\geq\frac{1}{2qd}n$. To complete the proof, we recall from Lemma 3.6 that
$\sigma_{q}(2n)\geq\sigma_{q}(2)\sigma_{q^{2}}(n)\geq\frac{n}{24qd}.$
@@
Lemma 3.5 implies $\sigma_{q}(l)\geq\sigma_{q}(2n)\geq n/24qd>l/96qd^{2}$ and this completes the proof. ∎
4 Analytic rank and Geometric rank under field extension
@@
This section is devoted to a discussion of the behaviour of $\operatorname{AR}_{\mathbb{K}}(T)$ and $\operatorname{GR}_{\mathbb{K}}(T)$ under field extensions, defined respectively in (6) and (4). For each $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and a field extension $\mathbb{F}/\mathbb{F}_{q}$, we denote by $T^{\mathbb{F}}$ the tensor in $\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}$ determined by $T$. To be more specific, we regard $T$ (resp. $T^{\mathbb{F}}$) as an $\mathbb{F}_{q}$-multilinear function (resp. $\mathbb{F}$-multilinear function) on $(\mathbb{F}_{q}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}_{q}^{n_{d}})^{\ast}$ (resp. $(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d}})^{\ast}$). Then $T^{\mathbb{F}}$ is obtained by extending $T$ linearly over $\mathbb{F}$. For instance,
$T^{\mathbb{F}}\left(a_{1}x_{1},\dots,a_{d}x_{d}\right)\coloneqq a_{1}\cdots a_{d}T(x_{1},\dots,x_{d})$
where $a_{1},\dots,a_{d}\in\mathbb{F}$ and $x_{1}\in(\mathbb{F}_{q}^{n_{1}})^{\ast},\dots,x_{d}\in(\mathbb{F}_{q}^{n_{d}})^{\ast}$.
We also recall that there is a canonical $\mathbb{F}$-isomorphism
$\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}\simeq\operatorname{Hom}_{\mathbb{F}}((\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast},\mathbb{F}^{n_{d}}),$
where $\operatorname{Hom}_{\mathbb{F}}((\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast},\mathbb{F}^{n_{d}})$ is the vector space consisting of all $\mathbb{F}$-multilinear maps from $(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast}$ to $\mathbb{F}^{n_{d}}$. We define the $\mathbb{F}_{q}$-linear map:
$\iota:\operatorname{Hom}_{\mathbb{F}}((\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast},\mathbb{F}^{n_{d}})\to\operatorname{Hom}_{\mathbb{F}_{q}}((\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast},\mathbb{F}^{n_{d}}),\quad\iota(L)=L.$ (13)
As an $\mathbb{F}_{q}$-multilinear map, we define $S_{\mathbb{F}_{q}}\coloneqq\iota(S)$. In the context of algebraic complexity theory [10, Chapter 15.3], $T^{\mathbb{F}}$ and $S_{\mathbb{F}_{q}}$ are called the scalar extension and restriction of $T$ and $S$, respectively.
Remark 4.1.
As a digression, we briefly discuss an alternative definition of $S_{\mathbb{F}_{q}}$ in terms of the trace map. Suppose that $\mathbb{F}$ is a separable extension of $\mathbb{F}_{q}$. We denote by $\operatorname{tr}:\mathbb{F}\to\mathbb{F}_{q}$ a trace map such that $\operatorname{tr}(c)=c$ for any $c\in\mathbb{F}_{q}$. Such a map exists since $\mathbb{F}$ is a separable extension of $\mathbb{F}_{q}$ (cf. [13, Proof of Corollary 1]). Given an $\mathbb{F}$-tensor $S\in\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}$, we may regard it either as an $\mathbb{F}$-multilinear map $\widetilde{S}:(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast}\to\mathbb{F}^{n_{d}}$ or as an $\mathbb{F}$-multilinear function $\overline{S}:(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d}})^{\ast}\to\mathbb{F}$. Canonically, $\widetilde{S}$ and $\overline{S}$ are related by
$\langle z_{d},\widetilde{S}(z_{1},\dots,z_{d-1})\rangle=\overline{S}(z_{1},\dots,z_{d}),\quad(z_{1},\dots,z_{d})\in(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d}})^{\ast}.$
Throughout this paper, we do not distinguish between $S$, $\widetilde{S}$ and $\overline{S}$, and we use $S$ to denote all of them.
Accordingly, $S_{\mathbb{F}_{q}}$ can also defined as an $\mathbb{F}_{q}$-multilinear function
$S_{\mathbb{F}_{q}}:(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d}})^{\ast}\to\mathbb{F}_{q},\quad S_{\mathbb{F}_{q}}(z_{1},\dots,z_{d})\coloneqq\operatorname{tr}(S(z_{1},\dots,z_{d})).$
We need to verify that $\iota(S)={\operatorname{tr}}\circ S$, when both $S$ and $\iota(S)$ are regarded as multilinear functions. Since both sides are $\mathbb{F}_{q}$-multilinear, it is sufficient to verify the equality on an $\mathbb{F}_{q}$-basis of $(\mathbb{F}^{n_{1}})^{\ast}\times\cdots\times(\mathbb{F}^{n_{d-1}})^{\ast}$. For each $1\leq k\leq d$, let $\{z^{(k)}_{1},\dots,z^{(k)}_{n_{k}}\}$ be an $\mathbb{F}$-basis of $(\mathbb{F}^{n_{k}})^{\ast}$, and let $\{x^{(k)}_{1},\dots,x^{(k)}_{n_{k}}\}$ be the $\mathbb{F}$-basis of $\mathbb{F}^{n_{k}}$ dual to $\{z^{(k)}_{1},\dots,z^{(k)}_{n_{k}}\}$. Suppose that $\{\beta_{1},\dots,\beta_{r}\}$ is an $\mathbb{F}_{q}$-basis of $\mathbb{F}$. Therefore, $\{\beta_{i}x^{(d)}_{j}:1\leq i\leq r,\;1\leq j\leq n_{d}\}$ is an $\mathbb{F}_{q}$-basis of $\mathbb{F}^{n_{d}}$. Let $\{f_{ij}:1\leq i\leq r,\;1\leq j\leq n_{d}\}$ be the $\mathbb{F}_{q}$-basis of $(\mathbb{F}^{n_{d}})^{\ast}$ dual to $\{\beta_{i}x^{(d)}_{j}:1\leq i\leq r,\;1\leq j\leq n_{d}\}$. We suppose
$\iota(S)(\beta_{i_{1}}z^{(1)}_{j_{1}},\dots\beta_{i_{d-1}}z^{(d-1)}_{j_{d-1}})=\sum_{i_{d},j_{d}=1}^{r,n_{d}}C^{i_{1},\dots,i_{d}}_{j_{1},\dots,j_{d}}\left(\beta_{i}x_{j}^{(d)}\right)\in\mathbb{F}^{n_{d}},$
where $C^{i_{1},\dots,i_{d}}_{j_{1},\dots,j_{d}}\in\mathbb{F}_{q}$. By definition we also have
$S(\beta_{i_{1}}z^{(1)}_{j_{1}},\dots\beta_{i_{d-1}}z^{(d-1)}_{j_{d-1}})=\sum_{i_{d},j_{d}=1}^{r,n_{d}}\beta_{i_{d}}C^{i_{1},\dots,i_{d}}_{j_{1},\dots,j_{d}}x_{j_{d}}^{(d)}\in\mathbb{F}^{n_{d}}.$
Therefore, we obtain
@@
$\operatorname{tr}\left(\langle S(\beta_{i_{1}}z^{(1)}_{j_{1}},\dots\beta_{i_{d-1}}z^{(d-1)}_{j_{d-1}}),f_{ij}\rangle_{\mathbb{F}}\right)$ $=\sum_{i_{d},j_{d}=1}^{r,n_{d}}C^{i_{1},\dots,i_{d}}_{j_{1},\dots,j_{d}}\operatorname{tr}\left(\langle\beta_{i_{d}}x^{(d)}_{j_{d}},f_{ij}\rangle_{\mathbb{F}}\right)$
$=C^{i_{1},\dots,i_{d-1},i}_{j_{1},\dots,j_{d-1},j}$
$=\langle\iota(S)(\beta_{i_{1}}z^{(1)}_{j_{1}},\dots\beta_{i_{d-1}}z^{(d-1)}_{j_{d-1}}),f_{ij}\rangle_{\mathbb{F}_{q}}.$
Here $\langle x,f\rangle_{\mathbb{F}}$ (resp. $\langle x,f\rangle_{\mathbb{F}_{q}}$) denotes the natural paring of $x\in\mathbb{F}^{n_{d}}$ with $f\in(\mathbb{F}^{n_{d}})^{\ast}$ over $\mathbb{F}$ (resp. $\mathbb{F}_{q}$).
Let $M_{d,\mathbb{F}}\in\underbrace{\mathbb{F}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}}_{d\text{~copies}}$ be the tensor corresponding to the multiplication map $\varphi_{d}$ defined in (7). Suppose that $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$ is an $\mathbb{F}_{q}$-tensor. We recall from [10, Equation (15.14)] that
$(T^{\mathbb{F}})_{\mathbb{F}_{q}}=T\boxtimes_{\mathbb{F}_{q}}M_{d,\mathbb{F}}\in\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}^{n_{d}}.$
To simplify notation, we use $T^{\mathbb{F}}$ to denote both $T^{\mathbb{F}}$ and $(T^{\mathbb{F}})_{\mathbb{F}_{q}}$ when there is no risk of confusion.
@@
[…]
$=\sum_{i=1}^{d}n_{i}-n_{d}-\frac{\log_{|\mathbb{F}_{q}|}(\lvert Z_{d}(T\boxtimes_{\mathbb{F}_{q}}M_{d,\mathbb{F}})\rvert)}{l}.$
Since $\mathbb{F}_{q}^{n_{i}}\otimes_{\mathbb{F}_{q}}\mathbb{F}\simeq\mathbb{F}_{q}^{n_{i}l}$ as $\mathbb{F}_{q}$-vector spaces, we may regard $T\boxtimes_{\mathbb{F}_{q}}M_{d,\mathbb{F}}$ as a tensor in $\mathbb{F}_{q}^{n_{1}l}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}l}$. By Lemma 2.2 again, we have
$\log_{\lvert\mathbb{F}_{q}\rvert}\left(\lvert Z_{d}(T\boxtimes_{\mathbb{F}_{q}}M_{d,\mathbb{F}})\rvert\right)=\sum_{i=1}^{d}n_{i}l-n_{d}l-\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes_{\mathbb{F}_{q}}M_{d,\mathbb{F}}).$
@@
$\operatorname{AR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right)=\frac{\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes_{\mathbb{F}_{q}}M_{d,\mathbb{F}})}{l}.$ (14)
Denote $s\coloneqq\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}})$. By definition, there exist matrices $L_{i}\in\mathbb{F}_{q}^{s\times l}$ for $1\leq i\leq d$ such that $(L_{1}\otimes\cdots\otimes L_{d})M_{d,\mathbb{F}}=\operatorname{Id}_{s}$, where $\operatorname{Id}_{s}$ is the diagonal tensor in $\underbrace{\mathbb{F}_{q}^{s}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{s}}_{d\text{~copies}}$. Therefore, we have
$\left((I_{n_{1}}\boxtimes_{\mathbb{F}_{q}}L_{1})\otimes\cdots\otimes(I_{n_{d}}\boxtimes_{\mathbb{F}_{q}}L_{d})\right)(T\boxtimes_{\mathbb{F}_{q}}L_{d,\mathbb{F}})=T\boxtimes_{\mathbb{F}_{q}}\operatorname{Id}_{s}=T^{\oplus s}.$
@@
$\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes_{\mathbb{F}_{q}}M_{d,\mathbb{F}})\geq\operatorname{AR}_{\mathbb{F}_{q}}(T^{\oplus s})=s\operatorname{AR}_{\mathbb{F}_{q}}(T)\geq c_{d,q}l\operatorname{AR}_{\mathbb{F}_{q}}(T)$ (15)
@@
To conclude the proof, we replace $\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}})$ by $\operatorname{R}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}})$ in the above argument and the other inequality follows immediately from Theorem 3.2. ∎
@@
It is proved in [35, Theorem 8.1] that for any $T\in\mathbb{Z}^{n_{1}}\otimes\cdots\otimes\mathbb{Z}^{n_{d}}$,
$\liminf_{p}\operatorname{AR}_{\mathbb{F}_{p}}\left(T^{\mathbb{F}_{p}}\right)=\operatorname{GR}_{\mathbb{Z}}(T),$ (18)
@@
According to [1, Proof of Theorem 1.13 (1)], we have $C\leq(1-\log_{q}(d))^{-1}$. Furthermore, the proof of Theorem 4.2 indicates that $c\geq(96d^{2}q)^{-1}$. As far as we are aware, the existence of $c$ is not previously known. Almost at the same time as this paper, Proposition 4.4 was independently proved in [3, Theorem 1.2] by another method.
Let $\mathbb{F}$ be a field and let $\langle n,n,n\rangle_{\mathbb{F}}$ be the tensor in $\mathbb{F}^{n}\otimes\mathbb{F}^{n}\otimes\mathbb{F}^{n}$ corresponding to the $n\times n$ matrix multiplication. By [35, Theorem 3], we have $\operatorname{GR}_{\mathbb{F}}(\langle n,n,n\rangle)=\lceil 3n^{2}/4\rceil$. Applying Theorems 4.2, 4.3 and Proposition 4.4 to $\langle n,n,n\rangle_{\mathbb{F}_{q}}$, we obtain the following result, which may be of independent interest.
@@
5 Partition rank and slice rank under field extension
In this section, we investigate the behaviour of the partition rank and the slice rank under field extensions. Let $\mathbb{K}$ be a field and let $\mathbb{F}$ be a finite extension of $\mathbb{K}$. We recall that any $\mathbb{F}$-tensor $S\in\mathbb{F}^{n_{1}}\otimes_{\mathbb{F}}\cdots\otimes_{\mathbb{F}}\mathbb{F}^{n_{d}}$ restricts to a $\mathbb{K}$-tensor $S_{\mathbb{K}}\in\mathbb{F}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{F}^{n_{d}}$ defined by a map similar to that in (13). Thus, both $\operatorname{PR}_{\mathbb{F}}(S)$ and $\operatorname{PR}_{\mathbb{K}}(S)\coloneqq\operatorname{PR}_{\mathbb{K}}(S_{\mathbb{K}})$ are well-defined. Moreover, they are related by an inequality.